Grade 12Physical Chemistry

3 Gms of activated charcoal was added to 50 ml of acetic acid (0.06N) in a flask. After an hour it was filtered and strength of the filterate was found to be 0.042N. The amount of acetic acid adsorbed will be

Gandikota Poorna Chandra
7 Years agoGrade 12

1 Answer

Arun
7 Years ago

As we learnt in

Concept of limiting reagent and excess reagent -

The reactant which gets consumed and thus limits the amount of product formed is called the limiting reagent.

- wherein

 CH4 (g) + 2O(g) → CO2 (g) + 2 H2O (g) ;

In this reaction, if 1 mole of methane and 1 mole of oxygen are taken then oxygen would be limiting reagent.

 

 

 Number of equivalent adsorbed =50\times 0.06-50\times 0.042=0.9

Mass =\frac{0.9}{1000}\times 60\times 1000\ mg=54\ mg

Mass adsorbed per gm =\frac{54}{3}=18\ mg